OpenStudy (coolshubhs):

Show that the roots of the equation x^2+2(3a+5)x+2(9a^2+25)=0 are complex unless a=5/3.

OpenStudy (anonymous):

first step: Evaluate Discriminant

OpenStudy (anonymous):

and of course we know when discriminant of a quadratic is negative then roots are complex

OpenStudy (coolshubhs):

solving discriminant fully i get 12a(-3a+10)

OpenStudy (coolshubhs):

@Hero help

OpenStudy (anonymous):

check ur work again

OpenStudy (coolshubhs):

in the first step i get : (2(3a+5))^2-4(2(9a^2+5))

OpenStudy (coolshubhs):

should I multiply completely after this ?

hero (hero):

What @mukushla is telling you is accurate.

OpenStudy (coolshubhs):

well how's that?

OpenStudy (coolshubhs):

how can we say that..? and how can we say that it is complex unless only a=5/3

OpenStudy (coolshubhs):

yeah i know when discriminant is negative it is complex but how to show that? and that exception of a=5/3

OpenStudy (anonymous):

tnx hero just simplfy that expression completely and see what happens

OpenStudy (coolshubhs):

ok. lemme solve.

OpenStudy (coolshubhs):

-36a^2-100+120a

OpenStudy (anonymous):

oh sorry its ok

OpenStudy (coolshubhs):

-36a^2-100+120a

OpenStudy (coolshubhs):

yeah that's it what i get.

OpenStudy (anonymous):

\[-36a^2+120a-100=-4(9a^2-30a+25)\]

OpenStudy (coolshubhs):

yup!

OpenStudy (coolshubhs):

but it shouldn't mean that −4(9a2−30a+25) is negative.?!

OpenStudy (coolshubhs):

because this term " (9a2−30a+25)" may also be negative

OpenStudy (anonymous):

9a^2=(3a)^2 25=5^2

hero (hero):

I'm pretty sure that whole thing is supposed to be set equal to zero.

OpenStudy (coolshubhs):

how this? 9a^2=(3a)^2 25=5^2 ??

OpenStudy (coolshubhs):

@Hero if we set that equal to zero it would mean that roots are real and equal.

OpenStudy (coolshubhs):

because the discriminant would be zero.

OpenStudy (anonymous):

see coolshubhs\[-4(9a^2−30a+25)=-4(3a-5)^2\]this is negative unless\[-4(3a-5)^2=0\]or\[a=\frac{5}{3}\]

OpenStudy (coolshubhs):

yes .. you got this but i dont understand why you set this expression equal to zero?

OpenStudy (coolshubhs):

well thanks dude :)